OpenStudy (anonymous):

Bob is taking his son to look at colleges. The first college they plan to visit is 150 miles from their home. In the first hour they drive at a rate of 60 mph. If they want to reach their destination in 2 1/2 hours. what speed must they average for the remainder of their trip?

OpenStudy (jiteshmeghwal9):

for the first hour they drive at a speed of 60 mph which means the distance covered in first hour is 60 miles. remaining time =3/2 hr remaining distance=90 miles average speed=(total distance traveled/ total time taken) here total distance must be 90 miles as u have to find the average speed for remaining trip & total time will be 3/2 hr.

Directrix (directrix):

d = r*t 90 = r * 1.5 Solve for r @starwarsfan123

OpenStudy (dan815):

90 miles divided by 1 and 1/2 hour.

OpenStudy (dan815):

He travelled 60 miler for the first hour as given 60 mph. Then for the remaining 1 and a 1/2 hour he had yet to travel 90 miles to reach his total destination of 150 miles. So 90/(1.5) = 60 mph

OpenStudy (dan815):

oh lol i just realized there was a much simpler way of doign this quesiton LOLOL

OpenStudy (dan815):

so he went 60 for 1 then 90 for 1.5 and 90 = 1.5(60) so 1.5(60) for 1.5(1) 60 for 1 therefore he was already going at the right speed to meet his destination in time haha!

OpenStudy (dan815):

hello anyone here?